OpenStudy (anonymous):

if a ball is tossed from ground level upward at a speed of 44 m/s how long will it take to reach its maximum height

OpenStudy (anonymous):

H=V^2/(2*g)=44^2/(2*10)= 96.8 Not considering the air friction

OpenStudy (anonymous):

@ellina so 96 seconds?

OpenStudy (anonymous):

t=v/g=44/10=4.4s in 4.4 seconds

OpenStudy (anonymous):

@ellina what is that first equation for? sorry im confused

OpenStudy (anonymous):

nevermind

OpenStudy (anonymous):

@ellina where did you get 96.8 then 4.4 seconds

OpenStudy (anonymous):

you are asking the max height 96.8 meter is the height

OpenStudy (anonymous):

and it takes only 4.4 second to get that height

OpenStudy (anonymous):

OK , I didn't read the problems seriously,sorry

OpenStudy (anonymous):

@ellina the other part is .. how long will it take to return to ground level (the time for entire trip)

OpenStudy (anonymous):

then ,you will use the max height to caculate the rest time t=sqrt{2 times H div g}=sqrt{2times96.8div10}